# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ 2.1 Initial Approach
The pseudo-posteriors are simply obtained by treating $\widehat{\mathbf{M}}$ as a fixed design matrix for the $j$th surrogate regression and then applying Bayes’ rule. To obtain independent pseudo-posterior samples of $\Psi$, we first draw independent samples $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{ind}{\sim}\widetilde{\Pi}_{j}$ for $j=1,\ldots,p$. We then let $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top},\widetilde{L}=\widetilde{\Lambda}\widetilde{\Lambda}^{\top},\widetilde{\Sigma}=\mbox{diag}(\widetilde{\sigma}_{1}^{2},\ldots,\widetilde{\sigma}_{p}^{2}),$ and $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}$ denote the pseudo-posterior samples of $\Lambda,L=\Lambda\Lambda^{\top}$, $\Sigma$, and $\Psi$, respectively. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach obtains independent pseudo-posterior samples of the covariance matrix $\Psi$ in an embarrassingly parallel fashion, bypassing the need to carry out MCMC.
@@
A natural concern is whether different choices of $\widehat{\mathbf{M}}$ satisfying (7) would affect the distribution of $\widetilde{\Psi}$. The following result, with proof in Section A of the Supplementary Material, ensures that this is not the case.
@@
(i) For any $\widehat{\mathbf{M}}$ satisfying (7), we have $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}=n\mathbb{I}_{k}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}=nUU^{\top}.$ (ii) The pseudo-posterior distribution of $\widetilde{L},\widetilde{\Sigma},$ and $\widetilde{\Psi}$ obtained from the FABLE approach only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$.
In our experience, the pseudo-posterior samples $\widetilde{\Psi}$ perform well when estimating the true covariance matrix, both in terms of simulation studies and theoretical posterior contraction rates. However, we observed that the entry-wise pseudo-posterior credible intervals obtained from the samples $\widetilde{\Psi}$ tend to underestimate the uncertainty associated with estimating the entry-wise elements of the true covariance matrix, from the perspective of frequentist coverage. In Section 2.2, we build on the initial FABLE approach developed in this subsection and propose a coverage correction (CC) of FABLE called CC-FABLE. We provide provably accurate guarantees on posterior contraction rates and frequentist coverage of the CC-FABLE procedure in Section 3.
Lastly, we describe the heuristic leading to (7). From (4), the matrix $\mathbf{A}=\mathbf{Y}V/\sqrt{p}$ satisfies
$\mathbf{A}=\mathbf{M}\dfrac{\Lambda^{\top}V}{\sqrt{p}}+\dfrac{EV}{\sqrt{p}}.$
Let $C=V^{\top}\Lambda/\sqrt{p}\in\mathbb{R}^{k\times k}$. Based on the consistency of spectral estimates, we expect $E$ to be approximately independent of $V$ as both $n,p$ grow. As a result, we expect $EV/\sqrt{p}\approx 0$ for increasing $p$. This leads us to
$\mathbf{A}\approx\mathbf{M}C^{\top}$ (12)
or equivalently, $a_{i}\approx C\eta_{i}$ for $i=1,\ldots,n$, where $a_{i}$ is the $i$th row of $\mathbf{A}$ for $i=1,\ldots,n$. Since $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k}),$ the marginal density of $a_{i}$ is approximately given by $a_{i}\overset{iid}{\sim}N_{k}(0_{k},CC^{\top})$, for $i=1,\ldots,n$. This motivates the following estimator $\widehat{C}\widehat{C}^{\top}$ of $CC^{\top}$:
$\widehat{C}\widehat{C}^{\top}=\dfrac{1}{n}\mathbf{A}^{\top}\mathbf{A}=\dfrac{1}{np}D^{2},$
leading us to (6). Given any $\widehat{C}$ satisfying (6), we use (12) to propose $\widehat{\mathbf{M}}=\mathbf{A}(\widehat{C}^{\top})^{-1}$ as an estimate of $\mathbf{M}.$ In particular, the canonical choice $\widehat{\mathbf{M}}=\sqrt{n}U$ is the same as the estimator obtained by carrying out principal components analysis (PCA) on the matrix $\mathbf{Y}\mathbf{Y}^{\top}$ (Fan et al.,, 2023). However, (7) permits other choices corresponding to different choices of $\widehat{C}$ satisfying (6), providing a general framework for obtaining estimates of the latent factors.
@@ 2.2 Coverage Correction
Although the pseudo-posterior of $\widetilde{\Psi}$ obtained from (10) performed well in terms of point estimation across simulation studies, we observed under-coverage of the entry-wise credible intervals, particularly when the signal-to-noise ratios $R_{j}=\|\lambda_{j}\|_{2}^{2}/\sigma_{j}^{2}$ for $j=1,\ldots,p$ are large. To improve coverage, we propose a simple coverage correction (CC) of the FABLE approach, which we denote by CC-FABLE. We provide a theoretical justification of this approach in Section 3. We found that the correction leads to improved performance in simulation studies, particularly with increasing $R_{j}$s.
To improve coverage, we include an additional factor $\rho^{2}$ in the conditional pseudo-posterior variance when generating $\widetilde{\lambda}_{j}$ for $j=1,\ldots,p$ following (10). That is, we modify $\widetilde{\Pi}_{j}$ in (10) to
@@
${\Pi}_{j}^{C}\left(\lambda_{C,j},\sigma_{C,j}^{2}\right)=N_{k}\left(\lambda_{C,j}\middle|\mu_{j},\rho^{2}\sigma_{C,j}^{2}\mathbf{K}\right)\mbox{IG}\left(\sigma_{C,j}^{2}\middle|\dfrac{\gamma_{n}}{2},\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\right),$ (13)
for $j=1,\ldots,p$, where $\mathbf{K}=\{\mathbf{M}^{\top}\mathbf{M}+(\mathbb{I}_{k}/\tau^{2})\}^{-1}=\mathbb{I}_{k}/(n+\tau^{-2})$ and $\gamma_{n},\mu_{j},\delta_{j}^{2}$ are as in (11). We then proceed as before to draw coverage-corrected pseudo-posterior samples $(\lambda_{C,j},\sigma_{C,j}^{2})\overset{ind}{\sim}\Pi_{j}^{C}$ for $j=1,\ldots,p$. Letting $\Lambda_{C}=[\lambda_{C,1},\ldots,\lambda_{C,p}]^{\top}$ and $\Sigma_{C}=\mbox{diag}(\sigma_{C,1}^{2},\ldots,\sigma_{C,p}^{2})$, we construct a coverage-corrected (CC) pseudo-posterior sample of the covariance matrix as $\Psi_{C}=\Lambda_{C}\Lambda_{C}^{\top}+\Sigma_{C}.$ Having computed the hyperparameters $\gamma_{n}$ and $\delta_{j}^{2}$ for $j=1,\ldots,p$ as in (11), the pseudo-posterior mean $\widehat{\Psi}_{C}$ of $\Psi_{C}$ is available analytically, with its $uv$th entry for $1\leq u,v\leq p$ given by
$\begin{split}\widehat{\Psi}_{C,uv}&=\left\{\begin{array}[]{ll}\mu_{u}^{\top}\mu_{v},&\text{if $u\neq v,$}\\ \|\mu_{u}\|_{2}^{2}+\left(1+\dfrac{k\rho^{2}}{n+\tau^{-2}}\right)\dfrac{\gamma_{n}\delta_{u}^{2}}{\gamma_{n}-2},&\text{if $u=v,$}\end{array}\right.\end{split}$ (14)
provided $\gamma_{n}>2$. For finite $k,\rho^{2},$ and $\tau^{2}$, the pseudo-posterior mean $\widehat{\Psi}_{C}$ can approximated as $\widehat{\Psi}_{C}=G_{0}G_{0}^{\top}+\Delta,$ where $G_{0}=[\mu_{1},\ldots,\mu_{p}]^{\top}$ and $\Delta=\mbox{diag}(\delta_{1}^{2},\ldots,\delta_{p}^{2})$, as $n$ increases. From here on, we take $\widehat{\Psi}_{C}=G_{0}G_{0}^{\top}+\Delta$ to be the pseudo-posterior mean of $\Psi$ under CC-FABLE, unless mentioned otherwise.
We choose $\rho$ to ensure that a summary statistic of the entrywise asymptotic coverages is $1-\alpha$, such as their average or minimum. Let $B$ denote the $p\times p$ symmetric matrix with entries $b_{uv}$, where
$\begin{split}b_{uv}&=\left\{\begin{array}[]{ll}\left(1+\dfrac{\|\mu_{u}\|_{2}^{2}\|\mu_{v}\|_{2}^{2}+(\mu_{u}^{\top}\mu_{v})^{2}}{\mathcal{V}_{u}^{2}\|\mu_{v}\|^{2}+\mathcal{V}_{v}^{2}\|\mu_{u}\|_{2}^{2}}\right)^{1/2},&\text{if $u\neq v,$}\\ \left(1+\dfrac{\|\mu_{u}\|_{2}^{2}}{2\mathcal{V}_{u}^{2}}\right)^{1/2},&\text{if $u=v,$}\end{array}\right.\end{split}$ (15)
with $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$. For example, we can ensure that all the entrywise asymptotic coverages are at least $1-\alpha$ by letting $\rho=\sup_{u,v}b_{uv}$. Another possibility is to let $\rho$ be the root of a nonlinear equation, which ensures the average of the asymptotic entrywise coverages is $1-\alpha$; we discuss the associated nonlinear equation along with the rationale behind possible choices of $\rho$ at the end of Section 3.3. Based on our experiences with finite sample simulations, we recommend choosing $\rho$ to the mean of $b_{uv}$ for $1\leq u\leq v\leq p$ as a computationally efficient strategy. This performed well with both median and average entrywise coverages close to $1-\alpha$. In what follows, we only consider the CC-FABLE version of our approach for theoretical justification, simulation results, and the real data application.
@@ 2.3 Hyperparameter Choice
Before obtaining pseudo-posterior samples of the covariance matrix, one needs to carefully choose two key hyperparameters. These are the number of factors $k$ and the global variance of the factor loadings $\tau^{2}$. To estimate $k$, we use the approach described in Chen and Li, (2022). The authors construct an information criterion $\mbox{JIC}(k)$ by adding the penalty $\nu(k,n,p)=k(n\vee p)\log(n\wedge p)$ to the model deviance. The proposed $\mbox{JIC}(k)$ criterion is similar in spirit to the Akaike information criterion (AIC) or the Bayesian information criterion (BIC) (Akaike,, 1974; Schwarz,, 1978). Given an upper bound $\mathcal{K}_{0}$ for the possible values of $k$, we estimate $\widehat{k}=\min_{1\leq k\leq\mathcal{K}_{0}}\mbox{JIC}(k).$ To choose an upper bound $\mathcal{K}_{0}$ for plausible values of $k$, we consider the singular values $s_{1}\geq s_{2}\geq\ldots\geq s_{n\wedge p}$ of $\mathbf{Y}$ and let $\mathcal{K}_{0}$ be the smallest integer $\mathcal{K}$ such that $\left(\sum_{j=1}^{\mathcal{K}}s_{j}\right)/\left(\sum_{j=1}^{n\wedge p}s_{j}\right)\geq S_{0}$, where we typically let $S_{0}$ be $0.5$ or $0.75$. In our experience across simulation studies and applications to gene expression datasets, we found this choice of $\widehat{k}$ to perform well.

### v3

@@ 2.1 Overview
To obtain independent posterior samples of $\Psi$, we first draw independent samples $(\widetilde{\lambda}_{j},\widetilde{\sigma}_{j}^{2})\overset{ind}{\sim}\widetilde{\Pi}_{j}$ for $j=1,\ldots,p$. We then let $\widetilde{\Lambda}=[\widetilde{\lambda}_{1},\ldots,\widetilde{\lambda}_{p}]^{\top},\widetilde{L}=\widetilde{\Lambda}\widetilde{\Lambda}^{\top},\widetilde{\Sigma}=\mbox{diag}(\widetilde{\sigma}_{1}^{2},\ldots,\widetilde{\sigma}_{p}^{2}),$ and $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}$ denote the posterior samples of $\Lambda,L=\Lambda\Lambda^{\top}$, $\Sigma$, and $\Psi$, respectively. Our proposed Factor Analysis with BLEssing of dimensionality (FABLE) approach obtains independent posterior draws of the covariance matrix $\Psi$ in an embarrassingly parallel fashion, entirely bypassing the need to carry out MCMC. We refer to the posterior distribution arising from the sampling scheme in (10) as the FABLE-posterior throughout this paper.
The FABLE-posterior mean $\widehat{\Psi}$ is available explicitly, with its $(u,v)$th entry for $1\leq u,v\leq p$ given by
[…]
provided $\gamma_{n}>2$. For finite $k,\rho^{2},$ and $\tau^{2}$, the FABLE-posterior mean $\widehat{\Psi}$ can be approximated with $\widehat{\Psi}\approx G_{0}G_{0}^{\top}+\Delta,$ where $G_{0}=[\mu_{1},\ldots,\mu_{p}]^{\top}$, $\Delta=\mbox{diag}(\delta_{1}^{2},\ldots,\delta_{p}^{2})$, and accuracy improves as $n$ increases. For simplicity, we refer to $\widehat{\Psi}=G_{0}G_{0}^{\top}+\Delta$ as the FABLE-posterior mean of $\Psi$ in what follows.
2.2 Pre-estimating the Latent Factors
We now describe the heuristic leading to (7). From (4), the matrix $\mathbf{A}=\mathbf{Y}V/\sqrt{p}$ satisfies
$\mathbf{A}=\mathbf{M}\,\left(\dfrac{\Lambda^{\top}V}{\sqrt{p}}\right)+\dfrac{EV}{\sqrt{p}}=\mathbf{M}\,C^{\top}+\dfrac{EV}{\sqrt{p}},$
where $C=V^{\top}\Lambda/\sqrt{p}\in\mathbb{R}^{k\times k}$. Based on the consistency of spectral estimates (Chen et al.,, 2021), we expect $E$ to be approximately independent of $V$ as both $n,p$ grow. As a result, we expect $EV/\sqrt{p}\approx 0$ for increasing $p$. This leads us to
$\mathbf{A}\approx\mathbf{M}\,C^{\top}$ (15)
or equivalently, $a_{i}\approx C\eta_{i}$, where $a_{i}$ is the $i$th row of $\mathbf{A}$, for $i=1,\ldots,n$. Since $\eta_{i}\sim N_{k}(0,\mathbb{I}_{k}),$ the marginal density of $a_{i}$ is approximately given by $a_{i}\overset{iid}{\sim}N_{k}(0,CC^{\top})$, for $i=1,\ldots,n$. This motivates the following estimator $\widehat{C}\widehat{C}^{\top}$ of $CC^{\top}$:
$\widehat{C}\widehat{C}^{\top}=\dfrac{1}{n}\mathbf{A}^{\top}\mathbf{A}=\dfrac{D^{2}}{np},$
leading us to (6). Given any invertible $\widehat{C}$ satisfying (6), we use (15) to propose $\widehat{\mathbf{M}}=\mathbf{A}(\widehat{C}^{\top})^{-1}$ as an estimator of $\mathbf{M}.$ In particular, the canonical choice $\widehat{\mathbf{M}}=\sqrt{n}U$ obtained with $\widehat{C}=D/\sqrt{np}$ is the same as the estimator obtained by carrying out principal components analysis (PCA) on the matrix $\mathbf{Y}\mathbf{Y}^{\top}$ (Fan et al.,, 2024). However, (7) permits other choices corresponding to different choices of $\widehat{C}$ satisfying (6), providing a general framework for obtaining estimates of the latent factors.
@@ 2.2 Pre-estimating the Latent Factors
A natural concern is whether different choices of $\widehat{\mathbf{M}}$ satisfying (7) would affect the FABLE-posterior distribution of $\widetilde{\Psi}$. The following result, with proof in Section A of the Supplementary Material, ensures that this is not the case.
@@
(i) For any $\widehat{\mathbf{M}}$ satisfying (7), we have $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}=n\mathbb{I}_{k}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}=nUU^{\top}.$ (ii) The FABLE-posterior distribution of $\widetilde{L},\widetilde{\Sigma}$, and $\widetilde{\Psi}$ obtained from the FABLE approach only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$.
Unless mentioned otherwise, we let $\widehat{\mathbf{M}}=\sqrt{n}U$ from here on.
@@ 2.3 Coverage Correction
Based on numerous simulations, we observed that after sampling
@@
$\widetilde{\lambda}_{j}\mid\widetilde{\sigma}_{j}^{2}\sim N_{k}(\mu_{j},\widetilde{\sigma}_{j}^{2}\mathbf{K}),\quad\widetilde{\sigma}_{j}^{2}\sim\mbox{IG}(\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2),$
the credible intervals of the entries of $\widetilde{\Psi}$ usually underestimated the uncertainty associated with estimating the entry-wise elements of the true covariance matrix, under repeated sampling. To alleviate this issue, the FABLE algorithm introduces a coverage-correction coefficient $\rho^{2}\geq 1$ in the conditional variance when sampling the factor loadings $\widetilde{\lambda}_{j}$ conditional on $\widetilde{\sigma}_{j}^{2}$s, as described in (10). This coefficient $\rho^{2}$ is adaptively chosen to restore nominal average frequentist coverage over all the entries. We provide provably accurate guarantees on frequentist coverage of the proposed approach in Section 3.3.
We choose $\rho$ to ensure that a summary statistic, such as the average or minimum, of the $\binom{p}{2}$ entrywise asymptotic coverages equals $1-\alpha$ and describe the methodology as follows. First, for $1\leq u\leq v\leq p$, let $\widehat{q}_{uv}(\rho)$ represent the estimated asymptotic coverage as a function of $\rho$ for the $(u,v)$th entry of the true covariance when estimated by $\widetilde{\Psi}_{uv}=\widetilde{\lambda}_{u}^{\top}\widetilde{\lambda}_{v}+\sigma_{u}^{2}\,\mathbbm{1}(u=v)$. We derive the form of $\widehat{q}_{uv}(\rho)$ and show that it is monotonically increasing for $\rho\geq 1$ in Section 3.3. Next, suppose $B$ denotes the $p\times p$ symmetric matrix with entries $b_{uv}$, where
[…]
with $\mu_{j}=\sqrt{n}\,U^{\top}y^{(j)}/(n+\tau^{-2})$ and $\mathcal{V}_{j}^{2}=\|(\mathbb{I}_{n}-UU^{\top})y^{(j)}\|_{2}^{2}/n$ for $j=1,\ldots,p$. The $b_{uv}\geq 1$ are defined so that $\widehat{q}_{uv}(b_{uv})=1-\alpha.$ That is, $b_{uv}$ represents the coefficient for correcting the coverage of the $(u,v)$th entry of the true covariance with $\widetilde{\Psi}_{uv}$. As a consequence of the monotonicity of $\widehat{q}_{uv}(\rho)$ for $\rho\geq 1$, we can ensure that all the entrywise asymptotic coverages are at least $1-\alpha$ by letting $\rho=\sup_{u,v}b_{uv}$, as $\widehat{q}_{uv}(\sup_{j,j^{\prime}}b_{jj^{\prime}})\geq\widehat{q}_{uv}(b_{uv})=1-\alpha$ for any given $1\leq u\leq v\leq p.$ Alternatively, we can choose $\rho$ to ensure that the average coverage over all the $p(p+1)/2$ distinct entries of the covariance equals $1-\alpha$. This is obtained by solving $\rho$ such that
$\dfrac{1}{p(p+1)/2}\sum_{1\leq u\leq v\leq p}\widehat{q}_{uv}(\rho)=1-\alpha.$ (17)
Based on our experiences with finite sample simulations, we found $\bar{b}:=\sum_{1\leq u\leq v\leq p}b_{uv}/\{p(p+1)/2\}$ as a close approximation to the solution of (17). Thus, as a computationally efficient default choice, we recommend choosing $\rho=\bar{b}.$ This choice ensures that the average entrywise coverage is close to $1-\alpha$.
Before obtaining FABLE-posterior samples of the covariance matrix, one needs to carefully choose two key hyperparameters. These are the number of factors $k$ and the global variance of the factor loadings $\tau^{2}$. We now describe the procedure to select each hyperparameter.
@@ 2.4 Hyperparameter Choice
2.4.1 Tuning $k$
To estimate $k$, we implement the approach described in Chen and Li, (2022), based on minimizing a joint-likelihood-based information criterion $\mbox{JIC}(k)$ by adding the penalty $\nu(k,n,p)=k(n\vee p)\log(n\wedge p)$ to twice the fitted negative log-likelihood. Although their proposed $\mbox{JIC}(k)$ is similar in spirit to the Akaike information criterion (AIC) or the Bayesian information criterion (BIC) (Akaike,, 1974; Schwarz,, 1978), there are some key differences. The AIC/BIC criteria are routinely used in existing literature on rank selection in factor models based on marginal likelihood $y_{i}\mid\Lambda,\Sigma\overset{iid}{\sim}N_{p}(0,\Lambda\Lambda^{\top}+\Sigma),$ obtained by marginalizing out $\eta_{i}\overset{iid}{\sim}N_{k}(0,\mathbb{I}_{k})$ (Akaike,, 1987). However, such approaches scale poorly as both $n,p$ increase, primarily due to the inversion of the $p\times p$ covariance $\Psi=\Lambda\Lambda^{\top}+\Sigma$ for each candidate $k$ when evaluating the marginal likelihood. Instead, the JIC considers the joint-likelihood of $\Lambda,\mathbf{M},$ and $\Sigma$, where $\mathbf{M}$ is the $n\times k$ matrix of latent factors. Evaluating the joint-likelihood bypasses the inversion of $\Psi$ entirely. Unfortunately, as pointed out by Chen and Li, (2022), existing approaches such as AIC/BIC are unable to consistently estimate the true rank of the model from joint-likelihoods, as the number of parameters diverges with increasing sample size. The authors show that their proposed JIC approach alleviates such issues, ensuring both computational scalability and consistency of the estimate $\widehat{k}$.
We now discuss the approach to obtain the fitted log-likelihood. From (4), the joint-likelihood function for a given $k$ is
$\mathcal{L}(\mathbf{M}\Lambda^{\top},\Sigma)=\prod_{i=1}^{n}\prod_{j=1}^{p}\dfrac{1}{\sqrt{2\pi}\sigma_{j}}\exp\left\{-\dfrac{1}{2\sigma_{j}^{2}}(y_{ij}-\mu_{ij})^{2}\right\},$
where $\mu_{ij}$ is the $(i,j)$th entry of the signal matrix $\mathbf{M}\Lambda^{\top}.$ Although $\mathbf{M}\in\mathbb{R}^{n\times k}$ and $\Lambda\in\mathbb{R}^{p\times k}$ have different dimensions as $k$ changes, the signal matrix $\mathbf{M}\Lambda^{\top}\in\mathbb{R}^{n\times p}.$ Motivated from the SVD of $\mathbf{Y}$ for a given $k$ as in (5), we estimate $\mathbf{M}\Lambda^{\top}$ with $\widehat{\mathbf{M}\Lambda^{\top}}=UDV^{\top}$. Next, we obtain an estimate of the error variances $\Sigma$ by maximizing the profile likelihood
$\mathcal{L}\left(\widehat{\mathbf{M}\Lambda^{\top}},\Sigma\right)=\prod_{i=1}^{n}\prod_{j=1}^{p}\dfrac{1}{\sqrt{2\pi}\sigma_{j}}\exp\left\{-\dfrac{1}{2\sigma_{j}^{2}}(y_{ij}-\hat{\mu}_{ij})^{2}\right\}$
with respect to $\sigma_{1}^{2},\ldots,\sigma_{p}^{2}$, where $\hat{\mu}_{ij}$ is the $(i,j)$th entry of $\widehat{\mathbf{M}\Lambda^{\top}}=UDV^{\top}$. Maximizing the profile likelihood with respect to $\sigma_{1}^{2},\ldots,\sigma_{p}^{2}$, the estimate $\hat{\sigma}_{j}^{2}$ is obtained as $\hat{\sigma}_{j}^{2}=\sum_{i=1}^{n}(y_{ij}-\hat{\mu}_{ij})^{2}/n.$ With $\widehat{\Sigma}=\mbox{diag}(\hat{\sigma}_{1}^{2},\ldots,\hat{\sigma}_{p}^{2})$, the fitted log-likelihood is given by
$\mathcal{L}(\widehat{\mathbf{M}\Lambda^{\top}},\widehat{\Sigma})=\prod_{i=1}^{n}\prod_{j=1}^{p}\dfrac{1}{\sqrt{2\pi}\hat{\sigma}_{j}}\exp\left\{-\dfrac{1}{2\hat{\sigma}_{j}^{2}}(y_{ij}-\hat{\mu}_{ij})^{2}\right\}=\dfrac{(2\pi e)^{-np/2}}{\prod_{j=1}^{p}\hat{\sigma}_{j}^{n}}.$
Once we have calculated the fitted log-likelihood for a given $k$, we proceed to compute the JIC:
$\mbox{JIC}(k)=-2\log\mathcal{L}\left(\widehat{\mathbf{M}\Lambda^{\top}},\widehat{\Sigma}\right)+k(n\vee p)\log(n\wedge p).$
@@ 2.4.1 Tuning $k$
Given a plausible upper bound $\mathcal{K}_{0}$ for the possible values of $k$, we estimate $\widehat{k}=\min_{1\leq k\leq\mathcal{K}_{0}}\mbox{JIC}(k).$ To choose an upper bound $\mathcal{K}_{0}$ for plausible values of $k$, we consider the singular values $s_{1}\geq s_{2}\geq\ldots\geq s_{n\wedge p}$ of $\mathbf{Y}$ and let $\mathcal{K}_{0}$ be the smallest integer $\mathcal{K}$ such that $\left(\sum_{j=1}^{\mathcal{K}}s_{j}\right)/\left(\sum_{j=1}^{n\wedge p}s_{j}\right)\geq S_{0}$, where we recommend setting $S_{0}=0.95$. In our experience across simulation studies and applications to gene expression datasets, we found this choice of $\widehat{k}$ to perform well. We carry out simulation studies to assess the performance of $\widehat{k}$ in Section G of the Supplementary Material.
2.4.2 Tuning $\tau^{2}$